One number is 11 more than twice another number. If the sum of the numbers is twice their difference, find the numbers.
Answer by Boreal(15235)

smaller is x
larger is 2x+11
sum is 3x+11
difference is x+11
3x+11=2(x+11)
3x+11=2x+22
x=11
2x+11=33
sum is 44 and difference is 22
The numbers are 11 and 33.